# Towards Geometric-Photometric Joint Alignment for Facial Mesh Registration

### v1

@@ 6 Conclusion
We present an innovative geometric-photometric joint alignment approach for facial mesh registration through the utilization of differentiable rendering techniques, demonstrating robust performance under various facial expressions with occlusion changes. Our method addresses this challenging task by designing holistic rendering alignment with multiple cues, optimized using a multiscale regularized algorithm. Unlike previous methods, our semantic annotation-free approach does not require marker point tracking or a set of aligned meshes for training. It is fully automatic and efficiently executed on a consumer GPU with fast convergence. Experimental results show that our aligned meshes achieve high geometric accuracy, surpassing conventional ICP-based techniques amd the state-of-the-art method NPHM. We also validate the photometric alignment by comparing rendered images with captured multi-view images, demonstrating pixel-level alignment in key facial areas, including the eyes, mouth, nostrils, and even freckles.
Our method has some limitations. Firstly, we observed that certain moles and freckles are too small to impose significant constraints, leading to the deformation ceasing before achieving pixel-level alignment. Secondly, while the mouth-open template can be effectively registered to various expressions, there are instances where the presence of teeth and tongues occasionally misguides the mouth contours.
While our method currently focuses on static facial scans with a single mouth-open template, future extensions can consider multi-view video sequences. Incorporating temporal information can address skin tone changes during performance and improve deformation accuracy. Additionally, refining the rendering function to include more complex effects, and exploring advanced inverse rendering pipelines for reconstructing material properties simultaneously are potential avenues for future researches.

### v2

@@ 6 Conclusion
We propose an innovative geometric-photometric joint alignment approach for facial mesh registration through the utilization of differentiable rendering techniques, demonstrating robust performance under various facial expressions with visibility changes. Unlike previous methods, our semantic annotation-free approach does not require marker point tracking or pre-aligned meshes for training. It is fully automatic and can be executed on a consumer GPU. Experiments show that our method achieve high geometric accuracy, surpassing conventional ICP-based techniques and the state-of-the-art method NPHM. We also validate the photometric alignment by comparing rendered images with captured multiview images, demonstrating pixel-level alignment in key facial areas, including the eyes, mouth, nostrils, and even freckles.
Our method currently adopts a relatively simple lighting and reflection model, which limits its performance under varying lighting conditions and skin highlights. Additionally, features such as teeth and the tongue may occasionally lead to inaccuracies in aligning the mouth contours. Future improvements will focus on optimizing efficiency, enhancing the rendering function to simulate more complex effects, and implementing cross-subject alignment based on genus-0 properties.